# How to Leverage Digit Embeddings to Represent Numbers?

### v1

@@ Abstract
Apart from performing arithmetic operations, understanding numbers themselves is still a challenge for existing language models. Simple generalisations, such as solving 100+200 instead of 1+2, can substantially affect model performance Sivakumar and Moosavi (2023). Among various techniques, character-level embeddings of numbers have emerged as a promising approach to improve number representation. However, this method has limitations as it leaves the task of aggregating digit representations to the model, which lacks direct supervision for this process. In this paper, we explore the use of mathematical priors to compute aggregated digit embeddings and explicitly incorporate these aggregates into transformer models. This can be achieved either by adding a special token to the input embeddings or by introducing an additional loss function to enhance correct predictions. We evaluate the effectiveness of incorporating this explicit aggregation, analysing its strengths and shortcomings, and discuss future directions to better benefit from this approach. Our methods, while simple, are compatible with any pretrained model and require only a few lines of code, which we have made publicly available. [^1] https://github.com/jasivan/Number-Embeddings
@@ 1 Introduction
Numbers play an integral role in language Thawani et al. (2021), and they are crucial across various domains such as finance Chen et al. (2018), medicine Jullien et al. (2023) or even sarcasm Dubey et al. (2019). Despite, large language models improving their capacity in many tasks, numerical reasoning still poses a challenge (Hong et al., 2024). Recent advancements in enhancing numerical reasoning within language models have predominantly stemmed from using more extensive or higher-quality training datasets Li et al. (2022a); Yu et al. (2024), scaling up models Lewkowycz et al. (2022); Kojima et al. (2022), or integrating methods like chain-of-thought reasoning Wei et al. (2022b); Yue et al. (2024). The effectiveness of such methods is significantly amplified when applied in conjunction with larger model architectures. With smaller models, the improvement shown is often minimal, for example, Wei et al. (2022b) use of chain-of-thought on a 20B parameter model only showed a 2.5% improvement on the MAWPS Koncel-Kedziorski et al. (2016) dataset whereas it jumps to 14.7% with a 137B parameter model. In addition, many of these solutions are computationally expensive or inaccessible, alternatively we seek a low cost approach that may have minimal impact on small scale models but greater effects on larger models.
@@
One main problem for number understanding is that the widely used tokenisation methods, like Byte-Pair Encoding (BPE) Sennrich et al. (2016), work well for common words but not for numbers. Specifically, rarer numbers might be broken down into random and meaningless pieces. In light of this, digit tokenisation Spithourakis and Riedel (2018) stands out for its simplicity and efficacy at representing numbers. This technique involves breaking down numbers into their individual digits, reducing vocabulary size and ensuring all decimal numbers can be accurately represented enhancing numerical reasoning abilities across various model architectures, tasks, and datasets Geva et al. (2020); Petrak et al. (2023); Sivakumar and Moosavi (2023). However, the aggregation of digit embeddings into a complete number representation is implicitly handled by the model, which raises the question: can explicit aggregation using mathematical priors improve numerical understanding? In this paper, we investigate this hypothesis by integrating a mathematically grounded aggregation of digit embeddings explicitly, rather than relying solely on the model’s inherent capabilities. We propose a novel approach to number embedding that requires no changes to the model’s architecture or additional pretraining. Our hypothesis is that an effective aggregation should meet two criteria: (1) it should distinguish between distinct numbers, ensuring unique representations for each value, and (2) the aggregated embedding should reflect natural numerical proximity. We also explore two approaches for this integration: adding a special token before the representation of individual digits to enhance input number representations, and incorporating an additional loss function to improve the representation of output digits.
Our findings show that the integration of explicitly aggregated digit embeddings enhances performance on small-scale models, potentially leading to even greater improvements in larger models. The effectiveness of our integration strategy depends on the size and pretraining of the model used. Our proposed method has promising prospects thus we also enumerate some future directions to further improve number understanding, consequently numerical reasoning.
@@ 2 Related Work
Numerical reasoning is the ability to interact with numbers using fundamental mathematical properties and thus model an area of human cognitive thinking (Saxton et al., 2019). Given a maths worded problem, the model needs to interpret the relation between both numbers and the text to then solve the problem by means of arithmetic operations Ahn et al. (2024). Therefore, an accurate number representation is primordial to both distinguish between different numbers but also predict an accurate answer. The literature focuses on five different areas to better represent numbers.
@@ 2.1 Scaling
Increasing the number of parameters of pretrained models has improved their numerical reasoning but it is still nowhere near perfect. For example, Minerva (540B) Lewkowycz et al. (2022) continued to struggle with higher than seven digit multiplication. Moreover, Frieder et al. (2023) evaluate ChatGPT and GPT4 to conclude that these very large models are inconsistent in their response when answering mathematical questions ranging from arithmetic problems to symbolic maths. This suggest that the models lack fundamental understanding of maths and thus also numbers. One approach to improve number representation is to scale up the vocabulary by having more individual number tokens. For example, GPT3 has unique tokens from the numbers 0-520, whereas GPT4 has them up to 999. Despite general better performance of GPT4, it is not feasible to represent infinitely many numbers in finite memory capacity, making the vocabulary larger would increase the computational costs as well.
@@ 2.2 Tokenisation
A more practical approach for representing all numbers is digit tokenisation Spithourakis and Riedel (2018); Geva et al. (2020); this separates numbers into a sequence of individual digits. This method improves upon conventional wordpiece tokenisation as shown with GenBERT Geva et al. (2020) and Mistral-7B Jiang et al. (2023) by reducing vocabulary size and ensuring precise representation of all numbers. Despite its advantages over conventional tokenisation algorithms, digit tokenisation has limitations. It relies on the model to aggregate digit embeddings into complete number representations, a process for which the model lacks direct supervision. During pretraining, models typically learn to aggregate subword tokens effectively for common words. However, not all numbers are encountered frequently enough during pretraining for the model to learn accurate aggregation. As an example, when the same question is posed with numbers represented differently (once as an integer and once scaled to the thousands), FLAN large with digit tokenisation shows a performance drop of 10% Sivakumar and Moosavi (2023). This indicates that the model struggles with numerical consistency and accurate aggregation of digit embeddings.
@@ 2.3 Architectural level
Change in model architecture also aids numerical reasoning as shown by NumNET Ran et al. (2019) and xVAL Golkar et al. (2024). NumNET extracts the numbers from the input question and passage to create a directed graph with magnitude information about each number present, e.g. which is greater than the others. This information is passed to the model after encoding the input question to supplement it with comparative information about each number so that the model can use this to answer the query. Alternatively, xVAL generates two input encodings, one with the text where numbers are replaced by [NUM], and one with empty space for the text but the actual value of the number in their corresponding positions. From the number preserving encoding, each number is converted to vector embeddings that are composed of themselves at each entry. The product of this vector with the embedding of [NUM] is then injected into the first layer of the transformer for each number in the input sequence. For decoding, a bespoke process is created to extract the predicted number instead of outputting the [NUM] token. Despite the positive contributions of these papers, their methods lack versatility as they are not adaptable off-the-shelf to any pretrained model.
@@ 2.4 Loss Functions
Another approach to improve numerical reasoning is for models to intrinsically learn better representation by introducing an inductive bias in the loss function. A simple approach is Wallace et al. (2019)’s use of the mean squared error (MSE) loss across the batch to directly predict floats on a subset of DROP Dua et al. (2019) which consists of numerical answers. However, this method is limited to datasets that only predict numbers. Contrastive loss is also used to manipulate the representation of numbers, for instance, Petrak et al. (2023) draws nearer the representation generated by BPE and digit tokenisation of numbers through an auxiliary loss when doing extended pretraining to improve arithmetic reasoning in worded problems like DROP but also tables like SciGen Moosavi et al. (2021). Similarly, Li et al. (2022b) use contrastive learning but on computation trees. They first generate computation trees for the mathematical operations and use contrastive loss to pull nearer the graph representing the same operation, e.g. addition, and push other ones further. This is then integrated in the main loss and improves performance on two maths worded problem datasets, MathQA Amini et al. (2019) and Math23K Wang et al. (2017). While these loss functions are adaptable with different models, contrastive training is computationally expensive.
@@ 2.5 Input Representation
The most model agnostic method is changing the representation of the numbers in the input text. Wallace et al. (2019) explore worded forms of numbers, but this approach would overly rely on the tokeniser which would split them into subwords. Muffo et al. (2022) decomposes the numbers into place values in reverse order, e.g. 123 = 3 units, 2 tens, 1 hundreds which helps when working with remainders, e.g. when adding. However, this introduces many more tokens which is undesirable as well as either creating new vocabulary for each place value term or the danger of them being split into subword tokens. Zhang et al. (2020) preserves the numerical aspect and converts all numbers into scientific notation, e.g. 314.1 is represented as 3141[EXP]2, improving models’ ability to identify the magnitude of a number. Despite providing magnitudinal information, the number before [EXP] still needs to be represented. In fact, all the above strategies require the model to implicitly compute an overall aggregation for the numbers based on their individual components generated by the tokeniser of the model, whether these are digits or subwords. A simple, yet effective method is to introduce pause tokens before predicting the answer Goyal et al. (2024). This is evaluated by training a 1B parameter transformer model on C4 using [PAUSE] tokens and a 1% improvement is shown on the numerical reasoning dataset, GSM8K Cobbe et al. (2021). While this method can be used for inference only, they conclude that pretraining is recommended, therefore less applicable to existing models.
Our work is versatile within this line of research. Unlike previous methods that rely on the model to implicitly learn aggregation, we focus on the explicit aggregation of digit embeddings using mathematical priors. This provides direct supervision for the aggregation process, improving the accuracy of number representation. Furthermore, our method ensures that the embedding for a given number aligns with its numerical neighbours, enhancing the model’s numerical reasoning capabilities without altering the model architecture or requiring extensive retraining.
@@ 3 Aggregation of Digit Embeddings
We explore an approach which is a natural continuation of digit tokenisation as this has demonstrated its efficacy in enhancing numerical reasoning compared to BPE tokenisation. This improvement can be attributed to digit tokenisation’s utilisation of pretrained embeddings for individual digits, allowing the model to learn the overall representation through contextualised embeddings. In contrast, BPE may fragment longer and less frequent numbers into random subsequences, resulting in less meaningful aggregations than those achieved through digit tokenisation. However, the implicit aggregation process employed by digit tokenisation remains unclear; specifically, how the model forms the overall aggregation of a number given the embeddings of its individual digits.
In this paper, we investigate a mathematically motivated aggregation that takes into account the relative position of each digit within a number. Our approach generates an overall embedding for the number by considering the positional weight of each individual digit in that number. For example, given “123”, the common understanding of numbers as base-10 is “$1\times 100+2\times 10+3\times 1$”, so left most digits are weighted higher as they represent a greater portion of the number.
We design our weighted scheme such that (1) the embeddings of single-digit numbers remain intact, as these embeddings are effectively learned during pretraining, evidenced by the high performance of models on single-digit operations (Sivakumar and Moosavi, 2023), (2) the weights of consecutive place values increase exponentially to reflect base-10, and (3) the weights do not sum to 1, meaning that it is not normalising the sum, allowing for number composed of the same digits, e.g. “111” and “11”, to be represented differently. These properties would introduce a bias towards an accurate length of numbers and the correct digits from left to right as the left most digits are amplified, hence preserving natural numerical order.
@@
These weights are designed to satisfy three key properties. (1) Alignment with single-digit representations: when $N=1$, $w_{1}=1$, ensuring compatibility with the model’s pretraining on single digits. (2) Exponential growth: the exponential component $2^{N-i}$ mimics the base-10 system, providing an appropriate scale without causing the weights to grow too rapidly. This also ensures that the weights are not normalised. (3) Regularisation Term: the fractional component acts as a regularisation term, forming a normalised triangular number sequence. For instance, for a 3-digit number, the sequence is 1,3,6, normalised to 0.1,0.3,0.6. This ensures that the difference between consecutive digit weights increases proportionally, i.e., $w_{i}-w_{i-1}=w_{0}\times i$, replicating the exponential ratio between digit positions in a logarithmic space.
@@
We compare our bespoke weighted aggregation function to a more standard aggregation function, sum. For a set of digit embeddings, we apply these functions along each dimension to generate a unique embedding for the number represented by these digits. Figure 2 graphs the F1-score for both functions and different digit length, i.e. 2-digit would be the numbers 10 to 99. Appendix A has results for other aggregation functions: max, min, mean and median; these have the lowest alignment with natural order with an F1-score below 5%. These functions all have a normalising property meaning that the length of the number has no bearing on the aggregated embedding, as the functions only retrieve one entry for each dimension therefore cases like “$1$” would be equivalent to both “11” and “1”. Contrastingly, sum has better F1-scores for up to 3 digits as it possesses magnitudinal information since all the entries are summed up for each dimension distinguishing, for instance, a 2-digit set from a 3-digit set as it simply adds more numbers. However, it is position agnostic - it assigns equal weight to all the digit irrespective of their relative positions. Therefore, the embeddings generated from permutations of the same digits will always be equivalent, e.g. “85” and “58”. Since larger digit numbers have more such permutations, the F1-score reduces as the number of digits increases. Using this metric, the best aggregation is our weighted sum, the average F1-score rounds to 69% for 2 digits onwards suggesting that our weighted sum is closer to the ideal depiction in Figure 1. Undoubtedly, 1-digit F1-score is better as these embeddings are generated from pretraining, but also because the weighted scheme ensures that they are separated from the other number embeddings.
Despite this weighted scheme aligning the number embeddings with their natural order, the weights generated by Equation 1 can become excessively large after a certain point. This behaviour is, however, attenuated by the regularisation term which maintains the high F1-score of 69% for, at least, up to 6-digit long numbers.

### v2

@@ Abstract
Within numerical reasoning, understanding numbers themselves is still a challenge for existing language models. Simple generalisations, such as solving 100+200 instead of 1+2, can substantially affect model performance Sivakumar and Moosavi (2023). Among various techniques, character-level embeddings of numbers have emerged as a promising approach to improve number representation. However, this method has limitations as it leaves the task of aggregating digit representations to the model, which lacks direct supervision for this process. In this paper, we explore the use of mathematical priors to compute aggregated digit embeddings and explicitly incorporate these aggregates into transformer models. This can be achieved either by adding a special token to the input embeddings or by introducing an additional loss function to enhance correct predictions. We evaluate the effectiveness of incorporating this explicit aggregation, analysing its strengths and shortcomings, and discuss future directions to better benefit from this approach. Our methods, while simple, are compatible with any pretrained model, easy to implement, and have been made publicly available. [^1] https://github.com/jasivan/Number-Embeddings
How to Leverage Digit Embeddings to Represent Numbers?
Jasivan Alex Sivakumar  and Nafise Sadat Moosavi Department of Computer Science University of Sheffield United Kingdom {jasivakumar1|n.s.moosavi}@sheffield.ac.uk
@@ 1 Introduction
Numbers play an integral role in language Thawani et al. (2021), and they are crucial across various domains such as finance Chen et al. (2018), medicine Jullien et al. (2023) or even sarcasm Dubey et al. (2019). Despite, large language models improving their capacity in many tasks, numerical reasoning still poses a challenge (Hong et al., 2024). Recent advancements in enhancing numerical reasoning within language models have predominantly stemmed from using more extensive or higher-quality training datasets Li et al. (2022a); Yu et al. (2024), scaling up models Lewkowycz et al. (2022); Kojima et al. (2022), or integrating prompt-based strategies such as chain-of-thought reasoning Wei et al. (2022b); Yue et al. (2024). The effectiveness of such methods is significantly amplified when applied in conjunction with larger model architectures. With smaller models, the improvement shown is often minimal, for example, Wei et al. (2022b) use of chain-of-thought on a 20B parameter model only showed a 2.5% improvement on the MAWPS Koncel-Kedziorski et al. (2016) dataset whereas it jumps to 14.7% with a 137B parameter model. In addition, many of these solutions are computationally expensive or inaccessible; we seek a low-cost approach that may have minimal impact on small-scale models but greater effects on larger models.
@@
A key challenge for number understanding is that widely used tokenisation methods, like Byte-Pair Encoding (BPE) Sennrich et al. (2016), work well for common words but not for numbers. Specifically, rarer numbers might be broken down into random and meaningless pieces. In light of this, digit tokenisation Spithourakis and Riedel (2018) stands out for its simplicity and efficacy at representing numbers. This technique involves breaking down numbers into their individual digits, reducing vocabulary size and ensuring all decimal numbers can be accurately represented enhancing numerical reasoning abilities across various model architectures, tasks, and datasets Geva et al. (2020); Petrak et al. (2023); Sivakumar and Moosavi (2023). However, the aggregation of digit embeddings into a complete number representation is implicitly handled by the model, which raises the question: Can explicit aggregation using mathematical priors improve numerical understanding?
In this paper, we investigate this question by integrating a mathematically grounded aggregation of digit embeddings explicitly, rather than relying solely on the model’s inherent capabilities. Our findings show that this aggregated digit embedding enhances performance on small-scale models by up to 6.17% compared to our baseline without it, potentially leading to even greater improvements in larger models. However, the effectiveness of our integration strategy depends on the size, the architecture of the model used, and how these priors are integrated in the model. Our main contributions are as follows:
  • We propose a novel approach to number embedding that requires no changes to the model’s architecture or additional pretraining, by showing that an aggregation is effective if it meets two criteria: (1) it distinguishes between distinct numbers, ensuring unique representations for each value, and (2) the aggregated embedding reflects natural numerical proximity.
  • We explore two approaches for integrating our aggregation: adding a special token before the representation of individual digits to enhance input number representations, and incorporating an additional loss function to improve the representation of outputted digits.
@@ 2 Related Work
Numerical reasoning is the ability to interact with numbers using fundamental mathematical properties and model an area of human cognitive thinking (Saxton et al., 2019). Given a maths worded problem, for example “Sarah has 5 apples and eats 2. How many apples does she have left?”, the model needs to interpret the relation between both the numbers and the text to then solve the problem by means of arithmetic operations Ahn et al. (2024). Therefore, an accurate number representation is primordial to distinguish between different numbers and also predict an accurate answer. The literature focuses on five different areas to better represent numbers.
@@ 2.1 Scaling
Increasing the number of parameters of pretrained models has improved their numerical reasoning but it is still nowhere near perfect. For example, Minerva (540B) Lewkowycz et al. (2022) continues to struggle with greater than seven digit multiplication. Moreover, Frieder et al. (2023) found that very large models like ChatGPT and GPT4 are inconsistent when answering mathematical questions ranging from arithmetic problems to symbolic maths. This suggests that the models lack fundamental understanding of numbers and thus also maths. One approach to improve number representation is to scale up the vocabulary by having more individual number tokens. For example, GPT3 Brown et al. (2020) has unique tokens from the numbers 0-520, whereas GPT4 OpenAI (2023) has them up to 999. Despite general better performance of GPT4, it is not feasible to represent infinitely many numbers in finite memory capacity. Making the vocabulary larger also increases the computational costs.
@@ 2.2 Tokenisation
A more practical approach for representing all numbers is digit tokenisation Spithourakis and Riedel (2018); Geva et al. (2020) which separates numbers into a sequence of individual digits. This method improves upon conventional wordpiece tokenisation as shown with GenBERT Geva et al. (2020) and Mistral-7B Jiang et al. (2023) by reducing vocabulary size and ensuring precise representation of all numbers. Despite its advantages over conventional tokenisation algorithms, digit tokenisation has limitations. It relies on the model to aggregate digit embeddings into complete number representations. During pretraining, models typically learn to aggregate subword tokens effectively for common words. However, not all numbers are encountered frequently enough during pretraining for the model to learn accurate aggregation. As an example, when the same question is posed with numbers represented differently (once as an integer and once scaled to the thousands), FLAN large with digit tokenisation shows a performance drop of 10% Sivakumar and Moosavi (2023). This indicates that the model struggles with numerical consistency and accurate aggregation of digit embeddings.
@@ 2.3 Architectural level
Change in model architecture also aids numerical reasoning as shown by NumNET Ran et al. (2019) and xVAL Golkar et al. (2024). NumNET extracts the numbers from the input question and passage to create a directed graph with relative magnitude information about each number present, e.g. which is greater than the others. After encoding the input question, this comparative information is also passed to the model to also be leveraged in answering the query. Alternatively, xVAL generates two input encodings, a text only encoding where numbers are replaced by [NUM], and a number encoding with empty space for the text but the actual value of the number in each number’s corresponding position. From the number preserving encoding, each number is converted to a vector with each entry as the number itself. The product of this vector with the embedding of [NUM] is then injected into the first layer of the transformer for each number in the input sequence. For decoding, a bespoke process is created to extract the predicted number instead of outputting the [NUM] token. Despite the positive contributions of these papers, their methods lack versatility as they are not adaptable off-the-shelf for any pretrained model.
@@ 2.4 Loss Functions
Another approach to improve numerical reasoning is for models to intrinsically learn better representation by introducing an inductive bias in the loss function. A simple approach is Wallace et al. (2019)’s use of the mean squared error (MSE) loss across the batch to directly predict floats on a subset of DROP Dua et al. (2019) which consists of numerical answers. Unfortunately, this method is limited to datasets that only predict numbers. Contrastive loss has also been used to manipulate the representation of numbers, for instance, Petrak et al. (2023) draws nearer the representation generated by BPE and digit tokenisation of numbers through an auxiliary loss when doing extended pretraining to improve arithmetic reasoning in worded problems like DROP but also tables like SciGen Moosavi et al. (2021). Similarly, Li et al. (2022b) use contrastive learning but on computation trees. They first generate computation trees for the mathematical operations. Then they use contrastive loss to pull nearer the graph representing the same operation, e.g. additions, and push other ones further. This is then integrated in the main loss and improves performance on two maths worded problem datasets, MathQA Amini et al. (2019) and Math23K Wang et al. (2017). While these loss functions are adaptable with different models, contrastive training is computationally expensive and requires annotated data.
@@ 2.5 Input Representation
The most model agnostic method involves changing the representation of the numbers in the input text. Wallace et al. (2019) explore worded forms of numbers, but this approach overly relies on the tokeniser which splits them into subwords. Muffo et al. (2022) decompose numbers into place values in reverse order, e.g. 123 = 3 units, 2 tens, 1 hundreds which helps when working with carry-on, e.g. when adding. However, this introduces many more tokens in the input which is undesirable, and also requires new vocabulary for each place value name. Zhang et al. (2020) convert all numbers into scientific notation, e.g. 314.1 is represented as 3141[EXP]2, improving models’ ability to identify the magnitude of a number. Despite providing magnitudinal information, the number before [EXP] still needs to be represented faithfully. In fact, all the above strategies require the model to implicitly compute an overall aggregation for the numbers based on their individual components generated by the tokeniser of the model, whether these are digits or subwords. A simple, yet effective method is to introduce pause tokens before predicting the answer Goyal et al. (2024). This is evaluated by training a 1B parameter transformer model on C4 using [PAUSE] tokens and a 1% improvement is shown on the numerical reasoning dataset, GSM8K Cobbe et al. (2021). While this method can be used for inference only, they conclude that pretraining is recommended, therefore less applicable to existing models.
Within this line of research, our work is more versatile. Unlike previous methods that rely on the model to implicitly learn aggregation, we focus on the explicit aggregation of digit embeddings using mathematical priors. This provides direct supervision for the aggregation process, improving the accuracy of number representation. Furthermore, our method ensures that the embedding for a given number aligns with its numerical neighbours, enhancing the model’s numerical reasoning capabilities without altering the model architecture or requiring extensive retraining.
@@ 3 Aggregation of Digit Embeddings
Digit tokenisation has demonstrated its efficacy in enhancing numerical reasoning compared to BPE tokenisation. This improvement can be attributed to digit tokenisation’s utilisation of pretrained embeddings for individual digits, allowing the model to learn the overall representation through contextualised embeddings. In contrast, BPE tends to fragment longer and less frequent numbers into random subsequences, resulting in less meaningful aggregations than those achieved through digit tokenisation. However, the implicit aggregation process employed by digit tokenisation remains unclear; specifically, how the model forms the overall aggregation of a number given the embeddings of its individual digits.
In this paper, we investigate a natural continuation of digit tokenisation, a mathematically motivated aggregation that takes into account the relative position of each digit within a number. Our approach generates an overall embedding for the number by considering the positional weight of each individual digit in that number. For example, given “123”, the common understanding of numbers as base-10 is “$1\times 100+2\times 10+3\times 1$”, so the left most digits are weighted higher as they have a larger impact on the value of the number.
We design our weighted scheme such that: (1) the embeddings of single-digit numbers remain intact, as these embeddings are effectively learned during pretraining, evidenced by the high performance of models on single-digit operations (Sivakumar and Moosavi, 2023), (2) the weights of consecutive place values increase exponentially to reflect base-10, and (3) the weights do not sum to 1, meaning that it does not normalise the sum, allowing for numbers composed of the same digits, e.g. “111” and “11”, to be represented differently. These properties would introduce a bias towards an accurate length of numbers and the correct digits from left to right as the left most digits are amplified, hence preserving natural numerical order.
@@
These weights are designed to satisfy three key properties. (1) Alignment with single-digit representations: when $N=1$, $w_{1}=1$, ensuring compatibility with the model’s pretraining on single digits. (2) Exponential growth: the exponential component $2^{N-i}$ mimics the base-10 system, providing an appropriate scale without causing the weights to grow too rapidly. This also ensures that the weights are not normalised. (3) Regularisation Term: the fractional component acts as a regularisation term, forming a normalised triangular number sequence. For instance, for a 3-digit number, the triangular sequence is 1,3,6, normalised to 0.1,0.3,0.6. This ensures that the difference between consecutive digit weights increases proportionally, i.e., $w_{i}-w_{i-1}=w_{0}\times i$, replicating the exponential ratio between digit positions in a logarithmic space.
@@
We compare our bespoke weighted aggregation function to more standard aggregation functions such as sum. For a set of digit embeddings, we apply these functions along each dimension to generate a unique embedding for the number represented by these digits. Figure 2 graphs the F1-score for our weighted function and sum over different digit lengths, i.e. 2-digit would be the numbers 10 to 99. Appendix A has the results for other aggregation functions: max, min, mean and median; these have the lowest alignment with natural order with an F1-score below 5%. These functions all have a normalising property meaning that the length of the number has no bearing on the aggregated embedding, as the functions only retrieve one entry for each dimension therefore cases like “$1$” would be equivalent to both “11” and “1”. Contrastingly, sum has better F1-scores for up to 3 digits as it possesses magnitudinal information since all the entries are summed up for each dimension distinguishing, for instance, a 2-digit set from a 3-digit set as it simply adds more numbers. However, it is position agnostic - it assigns equal weight to all the digits irrespective of their relative positions. Therefore, the embeddings generated from permutations of the same digits will always be equivalent, e.g. “85” and “58”. Since larger digit numbers have more such permutations, the F1-score reduces as the number of digits increases. Using this metric, the best aggregation is our weighted sum, the average F1-score rounds to 69% for 2 digits onwards suggesting that our weighted sum is closer to the ideal depiction in Figure 1. Undoubtedly, 1-digit F1-score is better as these embeddings are generated from pretraining, but also because the weighted scheme ensures that they are separated from the other number embeddings.
Despite this weighted scheme aligning the number embeddings with their natural order, the weights generated by Equation 1 can surpass the precision used making it too large after a certain point. However, this behaviour is attenuated by the regularisation term which maintains the high F1-score of 69% for, at least, up to 6-digit long numbers as shown in Figure 2. Theoretically, the newly formed number representation should contain numbers that beginning with the same digits and only vary at the unit level. For example, the neighbourhood of 4523 should contain all numbers of the form 452X where X is a digit from 0 to 9, therefore eight of these coincide in the 10-nearest neighbour, namely 4520, 4521, 4522, 4524, 4525, 4526, 4527, and 4528.